OpenStudy (anonymous):

f(x)=3x^4+9x^2+6 what are the zeros

OpenStudy (marissalovescats):

First think you want to do is factor out a 3

OpenStudy (marissalovescats):



OpenStudy (marissalovescats):

Then you will factor the inside You can check that by x^2 * x^2 = x^4 2x^2+1x^2=3x^2 2 * 1 =2